Nitric acid, HNO3 is a coourless, corrosive liquid used in thye manufacturing of Nitrogen fertilizers and explosives. In an experiment to develop new explosives for mining operations, a 28.5 g sample of HNO3 was poured into a beaker. How many moles of HNO3 are there in this sample of HNO3?

Grade:11

1 Answers

Kurma Anudeep
11 years ago

1 MOLE = 1+14+48=63g

x moles=28.5g

x=28.5/63

x=0.46 moles